A wheel rotatung at 1p.5rads-1 takes 5.2 to stop calculate



Angular acceleration



The number if revolution the wheel makes before it stops

1
Expert's answer
2021-11-22T10:12:25-0500

The angular acceleration:


"\alpha=\omega/t=1.5/5.2=0.29\text{ rad/s}^2."

The number of revolutions:


"\phi=\alpha t^2/2=3.92\text{ rad}."